OpenStudy (anonymous):

Prove that cot^2(x)/(cscx+1)=(1-sinx)/sinx

OpenStudy (cwrw238):

try putting csc x = 1/sin x and cot^2 x = ( 1 - sin^2 x)/ sin^2 x

OpenStudy (cwrw238):

- and see what it simplifies to

OpenStudy (anonymous):

\[(cosec ^{2}(x)-1)\div (cosec(x) +1)=cosecX-1= (1\div sinX)-1=(1-sinx)\div(sinX)\]

OpenStudy (anonymous):

Cwrw238, im sort of still confused:/ after you plug those in would you flip the 1/sinx to the top making it

OpenStudy (anonymous):